OpenStudy (anonymous):

What is the slope of the line that contains the points (-8, 7) and (-4, -5)? A) -3 B) 1/3 C) 3 D) -1/3

zarkam21 (zarkam21):

Are you familiar with the formula that solves these?

OpenStudy (anonymous):

\[slope of the line=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (rhystic):

whoops so slope is m = (y2-y1)/(x2-x1) or rise/run so (-5-7)/(-4-8) solve and you have your answer.

OpenStudy (anonymous):

@surjithayer @Rhystic Thank you, its been quite awhile since I've worked on problems like this, I know how to solve it now, Appreciate the help

OpenStudy (anonymous):

\[m=\frac{ -5-7 }{ -4-(-8) }\]

OpenStudy (anonymous):

@surjithayer I got 3/1 which is 3

OpenStudy (anonymous):

-12/4=-3